# Lower bounds for Seshadri constants on blow Ups of $\mathbb{P}^{2}$

### v2

@@ Abstract.
Let $\pi:X_{r}\rightarrow\mathbb{P}^{2}$ be a blow up of $\mathbb{P}^{2}$ at $r$ distinct points $p_{1},p_{2},\dots,p_{r}$. We study lower bounds for Seshadri constants of ample line bundles on $X_{r}$. First, we consider the case when $p_{1},p_{2},\dots,p_{r}$ are on a curve of degree $d\leq 3$. Then we assume that the points are very general and show that $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$ for any ample line bundle $L$ and any $x\in X_{r}$ if the Strong SHGH conjecture is true. We explore the relation between bounded negativity and Seshadri constants and study the Seshadri function on $X_{r}$.
@@
Seshadri constant, Bounded Negativity, Seshadri function
@@
Author was partially supported by a grant from Infosys Foundation
@@ 1. Introduction
Seshadri constants are used to study the local positivity of ample line bundles on a projective variety. They were introduced by Demailly in [3]. Consider a smooth complex projective variety $X$. The Seshadri constant of a nef line bundle $L$ on $X$ at a point $x\in X$ is defined as follows:
$\varepsilon(X,L,x)=\inf\bigg{\{}\frac{L\cdot C}{\text{mult}_{x}C}~{}~{}\bigg{|}~{}~{}C\text{ is a reduced irreducible curve in $X$ passing through $x$}\bigg{\}}.$
There is an alternate definition for the same, given as follows:
Consider the blow up $\pi_{x}:\widetilde{X}(x)\rightarrow X$ of $X$ at $x$. Let $E=\pi_{x}^{-1}(\{x\})$ denote the exceptional divisor. Then
$\varepsilon(X,L,x)=\sup\{t\in\mathbb{R}~{}~{}|~{}~{}\pi_{x}^{\ast}L-tE\text{ is a nef bundle on }\widetilde{X}(x)\}.$
This definition enables us to conclude that $\varepsilon(X,L,x)\leq\sqrt{L^{2}}$. This inequality motivates us to give following definition:
@@
Let $L$ be an ample line budle on a smooth projective surface $X$ and let $x\in X$. We say that $L$ is submaximal at $x$ if $\varepsilon(X,L,x)<\sqrt{L^{2}}$.
If $L$ is submaximal at $x$ then there exists a reduced and irreducible curve $C$ passing through $x$ such that $\varepsilon(X,L,x)=\frac{L\cdot C}{\text{mult}_{x}C}$. Such curves are called Seshadri curves of $L$ at $x$. If $\widetilde{C}$ is the strict transform of a Seshadri curve $C$ on $\widetilde{X}(x)$, then by the Hodge index theorem we have $\widetilde{C}^{2}<0$. A reduced irreducible curve $C$ on $X$ is said to be a $(-m)$-curve if $C^{2}=-m$ and $K_{X}\cdot C=m-2$, where $K_{X}$ is the canonical line bundle of $X$.
Further, we will define two more constants on a surface $X$ which are related to the Seshadri constant of an ample line bundle at a point. Firstly, we will define Seshadri constant of $X$ at point $x\in X$, denoted by $\varepsilon(X,x)$, as follows:
@@
Secondly, the Seshadri constant of a surface $X$ is defined to be:
$\varepsilon(X)=\inf_{x\in X}\varepsilon(X,x).$
More details about Seshadri constants and local positivity can be found in [16, Section 5] and [1]. One of the main problems in the area of Seshadri constants is to give lower bounds for $\varepsilon(X)$. Ein and Lazarsfeld proved that $\varepsilon(X,x)\geq 1$ for very general points $x\in X$; see [5, Theorem]. But Miranda showed that for every $\delta>0$ there is a surface $X$ and an ample line bundle $L$ on $X$ such that $\varepsilon(X,L,x)\leq\delta$ at a point $x\in X$ (see [5, Example 3.1] and [16, Example 5.2.1]). In this article, we will address this question when $X$ is a blow up of $\mathbb{P}^{2}$ at finitely many points.
Let $\pi:X_{r}\rightarrow\mathbb{P}^{2}$ be the blow up of $\mathbb{P}^{2}$ at $r$ distinct points $p_{1},p_{2},\dots,p_{r}$. Define $H$ to be the pullback of a line not passing through any $p_{i}$ and let $E_{i}=\pi^{-1}(\{p_{i}\})$ denote the exceptional divisor corresponding to $p_{i}$, for all $1\leq i\leq r$ Then the linear equivalence classes of $H,E_{1},E_{2},\dots,E_{r}$ form a basis for the class group $\text{Cl}(X_{r})$ of $X_{r}$. We have
@@
With this basis, the canonical bundle of $X_{r}$ is $K_{X_{r}}:=-3H+E_{1}+E_{2}+\dots+E_{r}$. As the Seshadri constant at point $x$ in $X_{r}$ has an alternative definition as defined above, we will often consider the blow up $\pi_{x}:\widetilde{X}_{r}(x)\to X_{r}$ at $x$. By abuse of notation, $E_{i}$ together with $E_{x}:=\pi_{x}^{-1}(\{x\})$ will denote the exceptional divisors for $\widetilde{X}_{r}(x)$.
For a smooth projective surface $X$, in a recent work, Schmidt [18] defined a function called the Seshadri function on $\text{Nef}(X)$,
$\varepsilon_{x}:\text{Nef}(X)\rightarrow\mathbb{R}$
defined as follows:
$\varepsilon_{x}(L)=\varepsilon(X,L,x).$
He proved that this function is continuous and is piecewise linear in a neighborhood of an ample line $L$ in $\text{Nef}(X_{r})$ with $\varepsilon(X,L,x)<\sqrt{L^{2}}$. He also showed that
$W_{x}=\{L\in\text{Nef}(X)~{}~{}|~{}~{}\varepsilon(X,L,x)<\sqrt{L^{2}}\}$
is nowhere dense when $X$ is an abelian surface [18, Lemma 4.1]. In last section we study the Seshadri function on $X_{r}$.
To study the Seshadri constant we will look at linear systems on $\mathbb{P}^{2}$ as well as on $X_{r}$. There are many conjectures proposed by several authors in this direction. We briefly recall these below.
Consider a linear system $\mathcal{L}=|dH-m_{1}E_{1}-m_{2}E_{2}-\dots-m_{r}E_{r}|$ for some non-negative integers $m_{1},m_{2},\dots,m_{r}$, and $d\geq 1$ on a blow up $X_{r}$ of $\mathbb{P}^{2}$ at very general points $p_{1},\dots,p_{r}$. We say that $\mathcal{L}$ is non-special if dim $\mathcal{L}$ is equal to the expected dimension of $\mathcal{L}$, denoted by edim $\mathcal{L}$ and defined as
edim =$\text{edim }\mathcal{L}=\max\Bigg{\{}{d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}-1,-1\Bigg{\}}.$
Following is one of the main conjectures in this area, independently proposed by Segre[19], Harbourne[12], Gimigliano[8] and Hirschowitz[15].
@@
The above version was given by Segre in 1964. There are several versions given by other authors also. It is known that SHGH is true when $r\leq 8$. We also recall the following conjecture which is stronger than the SHGH conjecture, known as Strong SHGH conjecture or Generalized SHGH conjecture [6, Conjecture 3.6]. This was proposed by Ł. Farnik, K. Hanumanthu, J. Huizenga D. Schmitz and T. Szemberg in 2020 in order to exhibit the existence of ample line bundles with an irrational Seshadri constant at every point of a general blow up of $\mathbb{P}^{2}$.
@@
(${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}\leq\max\Bigg{\{}{t+1\choose 2}-2,0\Bigg{\}}.$
@@
Let $X_{r}$ be a blow up of $r$ very general points of $\mathbb{P}^{2}$. If $C$ is an irreducible and reduced curve in $X_{r}$ such that $C^{2}<0$, then $C$ is a $(-1)$-curve.
@@
We say that a surface $X$ has bounded negativity if there exists an integer $b(X)$ such that $C^{2}\geq-b(X)$ for every irreducible and reduced curve $C$ on $X$. Over a positive characteristic field, there exists a surface without bounded negativity [14, Exercise V.1.10]. The following conjecture is open for surfaces over a field of characteristic zero.
Conjecture 1.5 (Bounded Negativity Conjecture).
For each smooth complex projective surface $X$ there exists an integer $b(X)>0$ such that
$C^{2}\geq-b(X)$
for every irreducible reduced curve $C\subset X$.
As mentioned above we are interested in lower bounds of $\varepsilon(X_{r})$. We give lower bounds when the blown up points are on a curve of degree $d\leq 3$ (see Theorem 2.1), when $r\leq 8$ (see Theorem 3.3 and Theorem 3.6). For arbitrary $r$, when the blown up points are very general, we show that $\varepsilon(X_{r})\geq 1$ assuming Conjecture 1.4 is true (see Theorem 4.2).
In Section 5, we study the relation between bounded negativity and Seshadri constants. We give a lower bound for Seshadri constants assuming Conjecture 1.5. In Section 6, we study the Seshadri function of $X_{r}$. Further, we discuss the submaximality of ample line bundles on $X_{r}$.
@@ 2. Blow ups of points on a curve
Let $X_{r}$ denote a blow up of $\mathbb{P}^{2}$ at $r$ points on a curve of degree $d$ in $\mathbb{P}^{2}$. In this section, we will exhibit a lower bound for the Seshadri constant of ample line bundles at any point of $X_{r}$.
@@
Let $p_{1},p_{2},\dots,p_{r}$ be r distinct points lying on a curve $C\subseteq\mathbb{P}^{2}$ of degree $d\leq 3$. Let $X_{r}$ be the denote the blow up of $\mathbb{P}^{2}$ at $p_{1},p_{2},\dots,p_{r}$. Then for any ample line bundle $L$ on $X_{r}$ and $x\in X_{r}$, we have
$\varepsilon(X_{r},L,x)\geq\begin{cases}1\text{ if $d=1,2$ or $d=3$ and $L\cdot(-K_{X_{r}})\geq 2$}\\ \frac{1}{2}\text{ if $d=3$ and $L\cdot(-K_{X_{r}})=1$}\end{cases}$
@@
By [12, Theorem III (a)], an ample line bundle on a smooth anticanonical rational surface $X$ is base point free if $L\cdot(-K_{X})\geq 2$. It is well-known that for an ample base point free line bundle $L$ on a surface $X$, $\varepsilon(X,L,x)\geq 1$ for every $x\in X$ [16, Example 5.1.18]. First, we consider the case $d=3$. If $L\cdot(-K_{X_{r}})\geq 2$ then $\varepsilon(X_{r},L,x)\geq 1$. If $L\cdot(-K_{X_{r}})=1$, $2L$ will be an ample base point free line bundle, which gives $\varepsilon(X_{r},2L,x)\geq 1$. This is equivalent to $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$.
To finish, assume that $d=1$ or $d=2$. Let $L$ be an ample line bundle on $X_{r}$ and let $\widetilde{C}$ be the strict transform of $C$ on $X_{r}$. Since $L$ is ample we have $L\cdot\widetilde{C}\geq 1$ and $L\cdot H\geq 1$. As $-K_{X_{r}}=\widetilde{C}+(3-d)H$ and $d\leq 2$, it follows that $L\cdot(-K_{X_{r}})\geq 2$. ∎
@@
Let $C$ be an irreducible and reduced curve of degree $d$ in $\mathbb{P}^{2}$ and $X_{r}$ the blow up of $r$ distinct smooth points $p_{1},p_{2},\dots,p_{r}\in C$. If $L=eH-n(E_{1}+\dots E_{r})$ is an ample line bundle on $X_{r}$ then,
$\varepsilon(X_{r},L)\geq\frac{1}{r}~{}\;~{}~{}\;\forall r\geq d^{2}+1.$
@@
Let  $\widetilde{C}=dH-E_{1}-\dots-E_{r}$ be the strict transform of $C$. As above, it is enough to show that $rL$ is globally generated. By [9, Theorem 3.2] for $r\geq d^{2}+1$, $rL$ is globally generated if $(re+3)d>r(rm+1)$. This is equivalent to $r(ed-rm)>r-3d$. This is true since $L\cdot\widetilde{C}=ed-rm\geq 1$. ∎
3. Seshadri constants on Blow ups of $r\leq 8$ distinct points

### v3

@@ Abstract.
Let $\pi:X_{r}\rightarrow\mathbb{P}^{2}$ be a blow up of $\mathbb{P}^{2}$ at $r$ distinct points $p_{1},p_{2},\dots,p_{r}$. We study lower bounds for Seshadri constants of ample line bundles on $X_{r}$. First, we consider the case when the points lie on a curve of degree $d\leq 3$, and the case when $r\leq 8$. We then assume that the points are very general and show that $\varepsilon(X_{r})\geq\frac{1}{2}$ if the Strong SHGH conjecture is true.
@@
Seshadri constants
@@ 1. Introduction
Throughout this article, we work over the field of complex numbers.
Seshadri constants are numerical invariants associated with an ample line bundle on a projective variety. They measure the local positivity of the line bundle and were introduced by Demailly in [3] to study the Fujita conjecture.
@@
Let $X$ be a smooth complex projective surface. The Seshadri constant of a nef line bundle $L$ on a surface $X$ at a point $x\in X$ is defined as:
[…]
We also define three related constants on a surface $X$, each connected to the Seshadri constant of an ample line bundle at a point. First, the Seshadri constant of $X$ at a point $x\in X$, denoted by $\varepsilon(X,x)$, is defined as follows:
@@
Second, for a fixed ample line bundle $L$, the Seshadri constant of $X$ at $L$ is defined as:
$\varepsilon(X,L)=\inf_{x\in X}\varepsilon(X,L,x).$
Finally, we define the Seshadri constant of $X$, $\varepsilon(X)$ as:
$\varepsilon(X)=\inf_{x\in X}\varepsilon(X,x)=\inf_{L\text{ ample}}\varepsilon(X,L).$
More details about Seshadri constants and local positivity can be found in [14, Section 5] and [1].
Computing exact values of Seshadri constants is difficult in general, and remains open even for blow-ups of $\mathbb{P}^{2}$. As a result, a central problem in this area is to determine effective lower bounds. A seminal result in this direction is due to Ein and Lazarsfeld [5], who proved that $\varepsilon(X,x)\geq 1$ for very general points $x\in X$. On the other hand, Miranda constructed examples where the Seshadri constant can be arbitrarily small (see [5, Example 3.1] and [14, Example 5.2.1]). These contrasting results motivate the following natural and fundamental question:
Question 1.2.
Does there exist a surface $X$ such that $\varepsilon(X)=0$?
In this article, we will address this question when $X$ is a blow up of $\mathbb{P}^{2}$ at finitely many points. Let $\pi:X_{r}\to\mathbb{P}^{2}$ be the blow up of $\mathbb{P}^{2}$ at distinct points $p_{1},p_{2},\dots,p_{r}$. We focus on obtaining lower bounds for Seshadri constants on $X_{r}$, under various configurations of $p_{1},p_{2},\dots,p_{r}$.
We obtain a lower bound for $X_{r}$ when $p_{1},p_{2},\dots,p_{r}$ lie on a curve of degree $d\leq 3$ as follows:
Theorem 3.1.
Let $p_{1},p_{2},\dots,p_{r}$ be r distinct points lying on a curve $C\subseteq\mathbb{P}^{2}$ of degree $d\leq 3$. Let $X_{r}$ denote the blow up of $\mathbb{P}^{2}$ at $p_{1},p_{2},\dots,p_{r}$. Then for all line bundle bundles $L$ and for all points $x\in X_{r}$ we have,
[…]
Then we study the case $r\leq 8$ and show that $\varepsilon(X_{r})\geq\frac{1}{2}$. More precisely, by Theorem 4.3 and Theorem 4.6 we have the following:
Theorem.
For $r\leq 8$, let $X_{r}$ denote the blow up of $r$ distinct points on $\mathbb{P}^{2}$ and $L$ be an ample line bundle on $X_{r}$. If $\varepsilon(X_{r},L,x)<1$, then
  1. (1)
    $r=8$,
  2. (2)
  3. (3)
    $L=-K_{X_{8}}$, and
  4. (4)
    $\varepsilon(X_{8},L,x)=\frac{1}{2}$.
Now to address the same for $r\geq 9$, we assume a conjecture concerning linear systems of plane curve called Generalized SHGH or Strong SHGH (see Conjecture 2.3), which is a stronger version of the famous SHGH conjecture and also we assume the blown up points are very general. As stated above Miranda showed that Seshadri constants can be arbitrarily small. In fact, Miranda’s construction shows that for every integer $m\geq 2$, there is a blow up $X\to\mathbb{P}^{2}$ at $d^{2}$ points on a degree $d$ plane curve (the choice of $d$ depends on $m$), a particular ample line bundle $L$ on $X$ and a particular point $x\in X$ such that $\varepsilon(X,L,x)=\frac{1}{m}$. In contrast we show that if $X\to\mathbb{P}^{2}$ is a blow up of $\mathbb{P}^{2}$ at very general points, then Conjecture 2.3 implies that $\varepsilon(X,L,x)\geq\frac{1}{2}$ for every ample line bundle $L$ on $X$ and every point $x\in X$:
Theorem 5.2.
Assume that Conjecture 2.3 is true. Let $\pi:X_{r}\to\mathbb{P}^{2}$ be the blow up of $\mathbb{P}^{2}$ at very general points $p_{1},p_{2},\dots,p_{r}$. Then
$\varepsilon(X_{r})\geq\frac{1}{2}.$
Further, if $\varepsilon(X_{r},L,x)<1$ for some ample line bundle $L$ and for some $x\in X$, then $\varepsilon(X_{r},L,x)=\frac{1}{2}$.
Acknowledgments
It is a pleasure to thank my advisor Krishna Hanumanthu for a meticulous reading of the manuscript and for many helpful discussions, suggestions, and comments. I thank the referee for a careful reading of the paper and numerous suggestions, which improved the paper. The author was partially supported by a grant from Infosys Foundation.
2. Preliminaries
This section recalls the key definitions and conjectures necessary for our results. We begin with an alternative characterization of the Seshadri constant that will be used extensively throughout the paper.
Lemma 2.1.
Let $X$ be a smooth projective surface and $x\in X$ be a point on $X$. Let $\pi_{x}:\widetilde{X}(x)\rightarrow X$ be the blow up of $X$ at $x$. Let $E=\pi_{x}^{-1}(\{x\})$ denote the exceptional divisor. Then
$\varepsilon(X,L,x)=\sup\left\{t\in\mathbb{R}_{\geq 0}~~\Big{|}~~\pi_{x}^{\ast}L-tE\text{ is a nef line bundle on }\widetilde{X}(x)\right\}.$
This definition enables us to conclude that $\varepsilon(X,L,x)\leq\sqrt{L^{2}}$. If $\varepsilon(X,L,x)<\sqrt{L^{2}}$ then there exists a reduced and irreducible curve $C$ passing through $x$ such that $\varepsilon(X,L,x)=\frac{L\cdot C}{\text{mult}_{x}C}$. Such curves are called Seshadri curves of $L$ at $x$. If $\widetilde{C}$ is the strict transform of a Seshadri curve $C$ on $\widetilde{X}(x)$, then by the Hodge index theorem we have $\widetilde{C}^{2}<0$.
A reduced irreducible curve $C$ on $X$ is said to be a (m)(-m)-curve if $C^{2}=-m$ and $K_{X}\cdot C=m-2$, where $K_{X}$ is the canonical line bundle of $X$.
[…]
@@ 2. Preliminaries
With this basis, the canonical bundle of $X_{r}$ is $K_{X_{r}}:=-3H+E_{1}+E_{2}+\dots+E_{r}$. As the Seshadri constant at a point $x$ on $X_{r}$ has an alternative definition as defined above, we will often consider the blow up $\pi_{x}:\widetilde{X}_{r}(x)\to X_{r}$ at $x$. By abuse of notation, $E_{i}$ together with $E_{x}:=\pi_{x}^{-1}(\{x\})$ will denote the exceptional divisors for $\widetilde{X}_{r}(x)$.
To study the Seshadri constant it is essential to look at linear systems on $\mathbb{P}^{2}$ as well as on $X_{r}$. Consider the complete linear system $\mathcal{L}=|dH-m_{1}E_{1}-m_{2}E_{2}-\dots-m_{r}E_{r}|$ for some non-negative integers $m_{1},m_{2},\dots,m_{r}$, and $d\geq 1$ on a blow up $X_{r}$ of $\mathbb{P}^{2}$ at very general points $p_{1},\dots,p_{r}$. The expected dimension of $\mathcal{L}$ is denoted by edim $\mathcal{L}$ and defined as
$\text{edim }\mathcal{L}=\max\Bigg{\{}{d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}-1,-1\Bigg{\}}.$
It is easy to see that dim $\mathcal{L}\geq$ edim $\mathcal{L}$. We say that $\mathcal{L}$ is non-special if dim =edim $\mathcal{L}$$\text${edim $\mathcal{L}$ and special otherwise.
Segre-Harbourne-Gimigliano-Hirschowitz (SHGH) conjecture is one of the main conjectures in this direction. Even though there are several versions of this conjecture proposed by Segre[16], Harbourne[11], Gimigliano[8] and Hirschowitz[13], we only state the version proposed by Segre.
@@
It is known that the SHGH conjecture is true when $r\leq 8$. We recall the following conjecture known as the Strong SHGH conjecture or Generalized SHGH conjecture [6, Conjecture 3.6], which is stronger than the SHGH conjecture. This was proposed by Ł. Farnik, K. Hanumanthu, J. Huizenga, D. Schmitz and T. Szemberg in 2020 in order to exhibit the existence of ample line bundles with an irrational Seshadri constant at every point of a general blow up of $\mathbb{P}^{2}$.
@@
${d+2\choose 2}-\displaystyle\sum_{i=1}^{r}{m_{i}+1\choose 2}\leq\max\Bigg{\{}{t+1\choose 2}-2,0\Bigg{\}}.$
@@
Let $X_{r}$ be a blow up of $r$ very general points of $\mathbb{P}^{2}$. If $C$ is an irreducible and reduced curve on $X_{r}$ such that $C^{2}<0$, then $C$ is a (1)(-1)-curve.
@@ 3. Blow ups of points on a curve
Let $X_{r}$ denote a blow up of $\mathbb{P}^{2}$ at $r$ points on a curve of degree $d$ on $\mathbb{P}^{2}$. In this section, we will exhibit a lower bound for the Seshadri constant of ample line bundles at any point of $X_{r}$.
@@
Let $p_{1},p_{2},\dots,p_{r}$ be r distinct points lying on a curve $C\subseteq\mathbb{P}^{2}$ of degree $d\leq 3$. Let $X_{r}$ be the blow up of $\mathbb{P}^{2}$ at $p_{1},p_{2},\dots,p_{r}$. Then for ample line bundle $L$ and for all points $x\in X_{r}$ we have,
[…]
@@
By [12, Theorem III.1 (a)], an ample line bundle on a smooth anti-canonical rational surface $X$ is base point free if $L\cdot(-K_{X})\geq 2$. It is well-known that for an ample base point free line bundle $L$ on a surface $X$, $\varepsilon(X,L,x)\geq 1$ for every $x\in X$ [14, Example 5.1.18]. First, we consider the case $d=3$. If $L\cdot(-K_{X_{r}})\geq 2$ then $\varepsilon(X_{r},L,x)\geq 1$. If $L\cdot(-K_{X_{r}})=1$, $2L$ will be an ample base point free line bundle, which gives $\varepsilon(X_{r},2L,x)\geq 1$. This is equivalent to $\varepsilon(X_{r},L,x)\geq\frac{1}{2}$.
Now assume $d=1$ or $d=2$. Let $L$ be an ample line bundle on $X_{r}$ and let $\widetilde{C}$ be the strict transform of $C$ on $X_{r}$. Since $L$ is ample we have $L\cdot\widetilde{C}\geq 1$ and $L\cdot H\geq 1$. As $K_{X_{r}}=\widetilde{C}+(3-d)H$ and $d\leq 2$, it follows that $L\cdot(-K_{X_{r}})\geq 2$. ∎
@@
Let $C$ be a smooth curve of degree $d$ on $\mathbb{P}^{2}$ and $X_{r}$ the blow up of $r$ distinct smooth points $p_{1},p_{2},\dots,p_{r}\in C$. If $L=eH-n(E_{1}+\dots+E_{r})$ is an ample line bundle on $X_{r}$ then,
$\varepsilon(X_{r},L)\geq\frac{1}{r},~\;~~\;\forall r\geq d^{2}+1.$
@@
Let  $\widetilde{C}=dH-E_{1}-\dots-E_{r}$ be the strict transform of $C$. As above, it suffices to show that $rL$ is globally generated. By [9, Theorem 3.2] for $r\geq d^{2}+1$, $rL$ is globally generated if $(re+3)d>r(rn+1)$. This is equivalent to r(edrn)>r3dr(ed-rn)>r-3d. This is true since $L\cdot\widetilde{C}=ed-rn\geq 1$. ∎
4. Blow ups of $r\leq 8$ distinct points